I think this is linear equations in 2 variables....
3v+2d=33.07
6v+6d=79.38
v=?,d=?

Answer by checkley77(12844)

3v+2d=33.07 muktiply this equation by -2 & add
6v+6d=79.38
-6v-4d=-66.14
------------------
2d=13.24
d=13.24/2
d=6.62 ans.
3v+2*6.62=33.07
3v+13.24=33.07
3v=33.07-13.24
3v=19.83
v=19.83/3
v=6.61 ans.
Proof:
6*6.61+6*6.62=79.38
39.66+39.72=79.38
79.38=79.38
----------------------
v=6.61,d=6.62